# Unsupervised Contrastive Analysis for Salient Pattern Detection using Conditional Diffusion Models

### v1

@@ 5.1 Experimental Data
We conduct initial experiments on CelebA Liu et al. (2015), containing 202,599 facial images with diverse attributes. We create a subset focusing on subjects wearing eyeglasses and those without accessories, resulting in 15,353 images divided into two distinct classes: 1) Eyeglasses (EG): Images with the ’Eyeglasses’ attribute and no other accessory-related attributes, used solely for evaluation as target images; 2) No Eyeglasses (NEG): Images without the ’Eyeglasses’ attribute or any accessory-related attributes, used for training as background images. More information on dataset partitions and pre-processing is provided in Appendices B.1 and B.2.
Medical Images
@@
To evaluate our common features encoder on medical data, we focus on tumor detection in brain MRIs. For a comprehensive comparison, we utilize the IXI IXI dataset as a background reference for training, as done in previous studies Behrendt et al. (2024); Iqbal et al. (2023). Evaluation is conducted on the Multimodal Brain Tumor Segmentation Challenge 2021 (BraTS21) Menze et al. (2014); Bakas et al. (2017); Baid et al. (2021) and the multiple sclerosis dataset (MSLUB) Lesjak et al. (2018), featuring tumor and Multiple Sclerosis (MS) samples, respectively. Notably, only T2-weighted images are used from all datasets. More information on preprocessing and data partitioning is provided in Appendix B.3.
5.2 Experimental Setup
Implementation Details
We compare our approach with state-of-the-art diffusion-based methods on CelebA Liu et al. (2015), including Diff-AE Preechakul et al. (2022), SepVAE Louiset et al. (2023), and DDIM Dhariwal and Nichol (2021). To ensure fair comparison, we use the official implementation of these methods and train them on our dataset partitions. For MRI experiments, we solely utilize our common features encoder alongside DDPM Wyatt et al. (2022), $p$DDPM Behrendt et al. (2024), and $m$DDPM Iqbal et al. (2023). These methods adopt distinct noise and objective functions, indirectly comparable to DDIM in facial images. Notably, they use structured simplex noise instead of Gaussian noise, capturing MRI image frequency distribution better. Additionally, they employ an alternative objective function to directly minimize $l1$ error between input image $x_{0}$ and its reconstruction $\hat{x_{0}}$ (Equation (9)). During training, MRI volumes are processed slice-wise, uniformly sampled with replacement. At test time, full volume reconstruction is achieved by iterating over all slices [^2] Implementation code of our method will be publicly available..
Evaluation Pipeline
During inference, the salient pattern map is obtained by the absolute difference between the input image $x_{0}$ and the reconstructed image $\hat{x_{0}}$, where higher values indicate larger reconstruction errors. To enhance the quality of these maps, inspired by Behrendt et al. (2024), we apply a median filter with a kernel size of 5 for score smoothing, followed by a morphological erosion operation for 10 iterations. Subsequently, the map is binarized, and the threshold is fine-tuned on the validation set to maximize the average DICE score. Additionally, we calculate the average Area Under the Precision-Recall Curve (AUPRC) on the test set.
5.3 Results
@@
Table 1: Reconstruction quality of methods trained on CelebA (training set) and tested on CelebA (test set). NEG - No Eyeglasses, EG - Eyeglasses. $\left|\Delta\right|=\left|NEG-EG\right|$.
@@
Diff-AE (No $\textrm{x}_{T}$)  Preechakul et al. (2022) $0.6287$ $0.6079$ $0.0208$ 0.01040.0104{0.0104}0.0104 $0.0119$ $0.0015$ $0.0847$ 0.10410.1041{0.1041}0.1041 $0.0194$
SepVAE Louiset et al. (2023) $0.5796$ $0.4773$ $0.1023$ $0.0153$ $0.0268$ $0.0115$ $0.2912$ $0.3322$ $0.0410$
DDIM Dhariwal and Nichol (2021) 0.96900.9690{0.9690}0.9690 $0.6568$ 0.3122 $\underline{0.0004}$ $0.0095$ $0.0091$ 0.00770.0077{0.0077}0.0077 $0.1385$ 0.1308
@@
5.3.1 CelebA Dataset
Table 1 presents the reconstruction quality results of our method alongside Diff-AE Preechakul et al. (2022), SepVAE Louiset et al. (2023), and DDIM Dhariwal and Nichol (2021). These models were trained on the CelebA dataset using only background samples (NEG) and evaluated on both background (NEG) and target (EG) images. Evaluation metrics include SSIM Wang et al. (2004) ($\uparrow$), LPIPS Zhang et al. (2018) ($\downarrow$), and MSE ($\downarrow$). Our method (w/ RE) achieves the highest SSIM (0.9763) and lowest MSE (0.0003) and LPIPS (0.0059) when reconstructing background examples (NEG). While our method (w/o RE) and DDIM produce similar results, they fail to capture common information and high-level details compared to our method (w/ RE), as visually demonstrated in Figure 2. Additionally, ablation studies investigate the reconstruction quality when different $t\in[1,T]$ with $T=1000$ are chosen for corrupting the input image to predict its denoised version (Figure 10 in Appendix D). Empirically, we find that $t=250$ strikes a balance between image quality and common information preservation.
@@ 5.3.1 CelebA Dataset
Table 1 provides metrics for distinguishing between the two classes (NEG vs. EG), with the magnitude $\left|\Delta\right|=\left|NEG-EG\right|$ indicating distinctiveness. Histogram plots are available in Figure 7 in Appendix C. Diff-AE Preechakul et al. (2022) and SepVAE Louiset et al. (2023) exhibit overlapping distributions across metrics, while our method demonstrates well-separated distributions, yielding the highest ROC AUC score (Figure 8 in Appendix C).
@@
Varying the margin $\epsilon$ in Equation (7) can influence identification accuracy by affecting the separation of different classes within the latent space. We conduct an ablation study on $\epsilon$ to observe its impact on identification accuracy using a $k$-NN classifier with $k=3$. Results depicted in Figure 3 (left) show that the highest accuracy is achieved when $\epsilon$ is chosen from the ranges $[0.5,2.5]$ with the RE transformation and $[0.15,1.0]$ without it.
@@
We conduct a linear probing analysis to assess the encoder’s class-invariance (Definition 4.2). The objective is to evaluate the encoder’s capability to distinguish between EG and NEG samples. Subsequently, a logistic regression classifier was trained on the extracted image latent features. Results in Figure 3 show lower target accuracy with RE, suggesting stronger common information preservation. However, instance classification performs worse with RE, possibly due to built-in augmentations in contrastive learning. Implementation details and numerical results are available in Appendix D.2.
OriginalDiff-AESepVAEDDIMOurs(w/o RE)Ours(w/ RE)OriginalDiff-AESepVAEDDIMOurs(w/o RE)Ours(w/ RE)Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
$\epsilon$$\epsilon$Refer to captionRefer to caption
Figure 3: Left: Instance classification (accuracy in %) for different values of $\epsilon$. Right: Classification accuracy (NEG vs. EG) for different values of $\epsilon$.
Table 2: Comparison of the evaluated models on brain MRI datasets. The best results are highlighted in bold, and the second-best results are underlined. * denotes that the results were obtained by reproducing the method. The coloured values between brackets indicate the improvement w.r.t. the correspondent method. The reported results for all metrics correspond to the mean $\pm$ standard deviation across the different folds.
@@
DICE (%)\;$\uparrow$ AUPRC (%)\;$\uparrow$ DICE (%)\;$\uparrow$ AUPRC (%)\;$\uparrow$ $l1(1e-3)\downarrow$
AE  Baur et al. (2021) $32.87$±$\pm 1.25$ $31.07$±$\pm 1.75$ $7.10$±$\pm 0.68$ $5.58$±$\pm 0.26$ $30.55$±$\pm 0.27$
VAE Baur et al. (2021) $31.11$±$\pm 1.50$ $28.80$±$\pm 1.92$ $6.89$±$\pm 0.09$ $5.00$±$\pm 0.40$ $31.28$±$\pm 0.71$
SVAE Behrendt et al. (2022) $33.32$±$\pm 0.14$ $33.14$±$\pm 0.20$ $5.76$±$\pm 0.44$ $5.04$±$\pm 0.13$ $28.08$±$\pm 0.02$
DAE Kascenas et al. (2022) $37.05$±$\pm 1.42$ 44.9944.99{44.99}44.99±$\pm 1.72$ $3.56$±$\pm 0.91$ $5.35$±$\pm 0.45$ $10.12$±$\pm 0.26$
f-AnoGAN Schlegl et al. (2019) $24.16$±$\pm 2.94$ $22.05$±$\pm 3.05$ $4.18$±$\pm 1.18$ $4.01$±$\pm 0.90$ $45.30$±$\pm 2.98$
DDPM* Wyatt et al. (2022) $39.25$±$\pm 1.01$ $47.79$±$\pm 1.28$ $5.43$±$\pm 1.71$ $7.25$±$\pm 0.79$ $14.10$±$\pm 1.64$
pDDPM* Behrendt et al. (2024) $49.47$±$\pm 0.91$ $54.68$±$\pm 1.02$ $9.17$±$\pm 1.29$ $\underline{10.35}$±$\pm 0.80$ $11.31$±$\pm 0.91$
mDDPM* Iqbal et al. (2023) $\underline{51.77}$±$\pm 0.18$ $\underline{57.49}$±$\pm 0.41$ $5.03$±$\pm 0.37$ $9.47$±$\pm 0.02$ 7.287.28{7.28}7.28±$\pm 0.00$
Ours (Common Features Encoder $+$ DDPM) 39.25±$\pm 0.55$ (+0.00) 48.58±$\pm 1.06$ (+0.79) $5.39$±$\pm 1.13$ 7.29±$\pm 0.23$ (+0.04) 13.72±$\pm 0.56$ (+0.38)
Ours (Common Features Encoder $+$ pDDPM) $49.36$±$\pm 0.53$ $54.60$±$\pm 0.60$ 10.48±$\pm 2.19$ (+1.31) 10.2610.26{10.26}10.26±$\pm 0.67$ 11.11±$\pm 0.01$ (+0.20)
Ours (Common Features Encoder $+$ mDDPM) $51.69$±$\pm 0.17$ 57.58±$\pm 0.17$ (+0.09) 9.47±$\pm 0.18$ (+4.44) $9.44$±$\pm 0.48$ $7.47$±$\pm 0.00$
Identity-conditional sampling
To demonstrate the efficacy of our encoder in capturing background patterns from a query image $x_{0}\sim q(x_{0})$, we encode $x_{0}$ to derive its latent representation $z_{bg}=f_{\theta}(x_{0})$. Then, using the step-wise denoising process $p_{\theta}(x_{t-1}|x_{t},z_{bg})$, we generate new images sharing the common information of $x_{0}$. Examples are shown in Figure 9 in Appendix C. With our encoder (w/ RE), the generated images preserve most background features and facial pose. Notably, subjects’ eyes appear more natural and perceptible with RE, suggesting the encoder is class-invariant, focusing on common patterns between the query and generated images.
5.3.2 Brain MRI Datasets
Table 2 provides a comparison between different unsupervised salient pattern detection methods in brain MRI. Evaluation metrics include DICE score, AUPRC, and $l1$ reconstruction error on healthy data (IXI). We reproduced results for DDPM Wyatt et al. (2022), pDDPM Behrendt et al. (2024), and mDDPM Iqbal et al. (2023). Results for other methods are referenced from the work of Behrendt et al. Behrendt et al. (2024).
Reconstruction Results
As depicted in Table 2, incorporating our common features encoder into diffusion-based models results in slight improvements across all datasets. Particularly noteworthy is the enhancement observed over the base results for the IXI dataset when our encoder is utilized. This suggests that when reconstructing a background (healthy) sample, the generated image preserves most of the common patterns from the input image. Detailed reconstruction examples can be found in Figure 11 in Appendix D.
Salient Pattern Detection
Examples of saliency maps generated by various methods are illustrated in Figure 4. Notably, we observe reduced foreground noise in the maps generated by $m$DDPM compared to DDPM, particularly evident in regions representing healthy brain anatomy. Specifically, when comparing the maps of the image depicted in the bottom-right row, the combination of Ours+DDPM and Ours+$m$DDPM demonstrates a smaller reconstruction error.
[…]
Figure 4: Salient pattern maps generated by different methods. Lighter colors indicate larger anomalies.

### v3

@@ 5.1 Datasets
We first conduct proof-of-concept experiments on CelebA (Liu et al., 2015) dataset, containing 202,599 facial images with diverse attributes. We create a subset focusing on subjects wearing eyeglasses and those without accessories, resulting in 15,353 images divided into two distinct classes: 1) Eyeglasses (EG): Images with the ’Eyeglasses’ attribute and no other accessory-related attributes, used solely for evaluation as target images; 2) No Eyeglasses (NEG): Images without the ’Eyeglasses’ attribute or any accessory-related attributes, used for training as background images. Our target task with this dataset is to remove glasses from input images while preserving subject-common information. Facial images are chosen due to their ease of visual interpretation, as identity preservation can be readily assessed during reconstruction. More details on dataset partitions and preprocessing are provided in the supplementary material.
Brain MRIs
@@
We evaluate our approach in brain MRIs for the task of tumor detection and localization. For a comprehensive comparison, we utilize the IXI [^2] https://brain-development.org/ixi-dataset/ dataset as a background (healthy) reference for training, as in previous studies (Behrendt et al., 2024, 2025). Evaluation is conducted in the Multimodal Brain Tumor Segmentation Challenge 2021 (BraTS21) (Menze et al., 2014; Bakas et al., 2017; Baid et al., 2021) and the multiple sclerosis dataset (MSLUB) (Lesjak et al., 2018), featuring tumor and Multiple Sclerosis (MS) samples, respectively. Furthermore, we evaluate our model on the recently evaluation-only NOVA (Bercea et al., 2025) dataset, comprising 906 brain MRI scans spanning 281 rare and diagnostically diverse pathologies. Notably, only T2-weighted images are used from all datasets. More details on data preprocessing partitioning are provided in the supplementary material.
5.2 Implementation Details
As target-invariant encoder, we use a ResNet-50 backbone with a fully connected layer and a 128-dimensional output. For brain MRI experiments, we adapt the Spark2D (Tian et al., 2023) framework to incorporate the $\epsilon$-InfoNCE loss. Notably, the encoder is pre-trained in the healthy training set. At the inference phase, we fine-tune it along with the denoising network. For the diffusion model, structured simplex noise (Wyatt et al., 2022) is applied. During training, timesteps $t\in[1,1000]$ are sampled uniformly, while inference uses a fixed value of $t=500$. The denoising network follows a U-Net architecture. Training runs for up to 1600 epochs on NVIDIA A40 GPUs with Adam optimizer (learning rate of 1e-4, and a batch size of 32). Code is publicly available at https://github.com/CristianoPatricio/unsupervised-contrastive-cond-diff.
6 Results
@@
Table 1: Reconstruction quality of methods trained on CelebA (training set) and tested on CelebA (test set). NEG - No Eyeglasses, EG - Eyeglasses. $\left|\Delta\right|=\left|NEG-EG\right|$. RE - Random Erasing.
@@
Diff-AE (Preechakul et al., 2022) $0.6287$ $0.6079$ $0.0208$ 0.01040.0104{0.0104}0.0104 $0.0119$ $0.0015$ $0.0847$ 0.10410.1041{0.1041}0.1041 $0.0194$
SepVAE (Louiset et al., 2023) $0.5796$ $0.4773$ $0.1023$ $0.0153$ $0.0268$ $0.0115$ $0.2912$ $0.3322$ $0.0410$
DDIM (Dhariwal & Nichol, 2021) 0.96900.9690{0.9690}0.9690 $0.6568$ 0.3122 $\underline{0.0004}$ $0.0095$ $0.0091$ 0.00770.0077{0.0077}0.0077 $0.1385$ 0.1308
@@
We begin by validating the effectiveness of our method on the CelebA dataset. Subsequently, we evaluated its performance in anomaly detection and localization tasks, comparing it with established state-of-the-art UAD methods.
6.1 Proof-of-Concept with CelebA Dataset
Table 1 presents the reconstruction quality results of our method against baseline models (Preechakul et al., 2022; Louiset et al., 2023; Dhariwal & Nichol, 2021). These models were trained in the CelebA dataset using only background samples (NEG) and evaluated in both background (NEG) and target (EG) images. Evaluation metrics include SSIM ($\uparrow$), LPIPS ($\downarrow$), and MSE ($\downarrow$). Our method (w/ RE) achieves the highest SSIM (0.9763) and the lowest MSE (0.0003) and LPIPS (0.0059) when reconstructing background examples (NEG). Although our method (w/o RE) and DDIM produce similar results, they fail to capture common information and high-level details compared to our method (w/ RE), as visually demonstrated in Figure 2. Furthermore, the supplementary material includes ablation studies analyzing reconstruction quality for different corruption timesteps $t\in[1,1000]$ and various values of $\epsilon$ in $\epsilon$-InfoNCE.
@@ 6.1 Proof-of-Concept with CelebA Dataset
Table 1 provides metrics for distinguishing between the two classes (NEG vs. EG), with the magnitude $\left|\Delta\right|=\left|NEG-EG\right|$ indicating distinctiveness (histogram plots are available in the supplementary material). Diff-AE and SepVAE exhibit overlapping distributions across metrics, while our method demonstrates well-separated distributions, yielding the highest ROC AUC score (see Figure C.3 in the supplementary material).
@@
Varying the margin $\epsilon$ in Equation (7) can influence identification accuracy by affecting the separation of different classes within the latent space. We conduct an ablation study on $\epsilon$ to observe its impact on identification accuracy using a $k$-NN classifier with $k=3$. Results depicted in Figure 3 (right) show that the highest accuracy is achieved when $\epsilon$ is chosen from the ranges $[0.5,2.5]$ with the random erasing (RE) transformation and $[0.15,1.0]$ without it.
Figure 3: Left: Classification accuracy (NEG vs. EG) for different values of $\epsilon$. Right: Instance classification (accuracy in %) for different values of $\epsilon$.
@@
We conduct a linear probing analysis to assess the encoder’s class-invariance (Definition 4.2). The objective is to evaluate the encoder’s capability to distinguish between EG and NEG samples. Subsequently, a logistic regression classifier was trained on the extracted image latent features. Results in Figure 3 (left) show lower target accuracy with RE, suggesting stronger common information preservation. However, instance classification performs worse with RE, possibly due to built-in augmentations in contrastive learning.
6.2 Application on Brain MRI Datasets
Table 2 compares different UAD methods on three brain MRI datasets using evaluation metrics such as DICE score, AUPRC, and $\ell_{1}$ reconstruction error (calculated only on healthy IXI dataset). We reproduce results for DDPM (Wyatt et al., 2022), pDDPM (Behrendt et al., 2024), and cDDPM (Behrendt et al., 2025), while results for other methods are taken from Behrendt et al. (2024).
Table 2: Comparison of evaluated models in brain MRI datasets. Best results are shown in bold, and second-best results are underlined. * indicate results obtained by reproducing the method. All metrics are reported as mean $\pm$ standard deviation across 5 folds. Dice score and AUPRC are reported in percentage (%).
@@ 6.2 Application on Brain MRI Datasets
DICE $\uparrow$ AUPRC $\uparrow$ DICE $\uparrow$ AUPRC $\uparrow$ $\ell_{1}(1e-3)\downarrow$
AE (Baur et al., 2021) $32.87$±$\pm 1.25$ $31.07$±$\pm 1.75$ $7.10$±$\pm 0.68$ $5.58$±$\pm 0.26$ $30.55$±$\pm 0.27$
VAE (Baur et al., 2021) $31.11$±$\pm 1.50$ $28.80$±$\pm 1.92$ $6.89$±$\pm 0.09$ $5.00$±$\pm 0.40$ $31.28$±$\pm 0.71$
SVAE (Behrendt et al., 2022) $33.32$±$\pm 0.14$ $33.14$±$\pm 0.20$ $5.76$±$\pm 0.44$ $5.04$±$\pm 0.13$ $28.08$±$\pm 0.02$
DAE (Kascenas et al., 2022) $37.05$±$\pm 1.42$ 44.9944.99{44.99}44.99±$\pm 1.72$ $3.56$±$\pm 0.91$ $5.35$±$\pm 0.45$ $10.12$±$\pm 0.26$
f-AnoGAN (Schlegl et al., 2019) $24.16$±$\pm 2.94$ $22.05$±$\pm 3.05$ $4.18$±$\pm 1.18$ $4.01$±$\pm 0.90$ $45.30$±$\pm 2.98$
DDPM* (Wyatt et al., 2022) $39.25$±$\pm 1.01$ $47.79$±$\pm 1.28$ $5.43$±$\pm 1.71$ $7.25$±$\pm 0.79$ $14.10$±$\pm 1.64$
pDDPM* (Behrendt et al., 2024) $49.47$±$\pm 0.91$ $54.68$±$\pm 1.02$ $9.17$±$\pm 1.29$ 10.35±$\pm 0.80$ $11.31$±$\pm 0.91$
cDDPM* (Behrendt et al., 2025) $50.57$±$\pm 1.34$ 56.12±$\pm 1.78$ $\underline{9.88}$±$\pm 0.83$ $9.45$±$\pm 0.60$ 9.71±$\pm 0.47$
Ours 50.97±$\pm 2.06$ $56.11$±$\pm 2.19$ 10.08±$\pm 1.52$ $\underline{9.65}$±$\pm 1.35$ $9.92$±$\pm 0.59$
Reconstruction Quality
To assess overall reconstruction quality, we use the held-out test set from the healthy IXI dataset and report the $\ell_{1}$-error, calculated as the average absolute difference between the reconstructed images and their corresponding inputs. As shown in Table 2, our approach achieves performance competitive with cDDPM and pDDPM, and significantly outperforms all other baselines. These results suggest that the model effectively preserves healthy anatomy from the input image during reconstruction. This can be attributed to the influence of the $\epsilon$-InfoNCE regularization (Eq. 7) within the target-invariant encoder, which improves the representation space with regard to the anatomy.
Tumor Segmentation
To evaluate segmentation performance in the context of UAD, we use all unhealthy test sets and report two metrics: i) Dice score and ii) Area Under Precision–Recall Curve (AUPRC). As shown in Table 2, our method is competitive with the top-performing baseline, showing marginal improvements in terms of Dice score. We achieve the best results with $\epsilon=0.5$ in Equation 7. Additional results with other $\epsilon$ values are available in the supplementary material. Analogous to the experimental findings on the CelebA dataset, this model can be used to distinguish between healthy and unhealthy MRIs by analyzing the generated saliency maps, as shown in Figure 1 (bottom-right). For instance, if the difference between a reconstructed brain MRI and its original counterpart shows high-intensity values across large regions, this may indicate presence of tumor and can be flagged for further review by a radiologist. This strategy can be seen as a form of clinical triage, while also contributing to the model’s explainability by providing a rationale for its decisions. Additional examples of reconstructions are shown in the supplementary material.
Anomaly Localization on NOVA Benchmark
Anomaly localization involves identifying and detect abnormal regions within brain MRIs. This task holds significant clinical importance, as many medical errors arise from failing to detect a pathology altogether. The objective is to predict one or more bounding boxes per image that correspond to abnormal regions, using annotations provided by radiologists as ground truth. As shown in Table 3, our method outperforms all baseline approaches in almost all evaluation metrics. Specifically, we achieve these results by considering only T2-weighted images and selecting the top-$5$ predicted largest bounding-boxes. We empirically found that including all modalities available in the NOVA dataset degrades performance, which is expected, as the models were pre-trained only on T2-weighted images from IXI. Compared to pDDPM and cDDPM, our pre-training strategy, which is based on contrastive learning, appears to be more robust than simply using data augmentation. As illustrated in Figure 4, our model is capable of detecting larger and more accurately localized anomalies, while baseline methods tend to produce smaller and more fragmented predictions. For example, pDDPM underestimates lesion size, likely due to its patching strategy, which may rely on suboptimal patch sizes. These results highlight the effectiveness of our methodology, particularly given that NOVA is considered a challenging and realistic evaluation-only benchmark dataset. Additional ablation studies on modality choice and the selection of top-$N$ largest bounding boxes are provided in the supplementary material.
Table 3: Localization performance on NOVA benchmark. Reported metrics include standard object detection measures (mean Average Precision at multiple IoU thresholds), detection accuracy (ACC50), number of true positives (TP30), and number of false positives (FP30).
Model mAP30 mAP50 mAP50-95 ACC50 TP30 FP30 $\downarrow$
DDPM (Wyatt et al., 2022) 17.50 8.23 2.75 7.36 63/394 1517
pDDPM (Behrendt et al., 2024) 0.30 0.30 0.06 0.25 1/394 1445
cDDPM (Behrendt et al., 2025) 11.51 5.19 1.51 4.72 44/394 1423
Ours 18.57 9.41 3.33 9.29 71/394 1492
Figure 4: Examples of model predictions for anomaly localization on NOVA. Compared to baseline models, our method produces more precise and spatially focused bounding boxes that better align with ground-truth annotations (green). Each row corresponds to a distinct pathology.